40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Jul 24 18:13:05 2013, Chloe, Notes, ECDL, Current Driver 
    Reply  Thu Jul 25 18:13:11 2013, Chloe, Notes, ECDL, Current Driver/TEC thermistor_temperature-resistance.xlsx
       Reply  Fri Jul 26 17:51:02 2013, Chloe, Notes, ECDL, Current Driver/TEC 
          Reply  Mon Jul 29 19:06:31 2013, Chloe, Notes, ECDL, Current Driver/TEC 
             Reply  Tue Jul 30 17:55:21 2013, Chloe, Notes, ECDL, Current Driver/TEC working!.jpg
Message ID: 1255     Entry time: Thu Jul 25 18:13:11 2013     In reply to: 1254     Reply to this: 1259
Author: Chloe 
Type: Notes 
Category: ECDL 
Subject: Current Driver/TEC 

Current driver: the plan for the summer is to use the current driver that Tara found, the LDX-3412. Dmass has a low noise current driver that he has developed that eventually (after the summer, since it is unfeasible to finish over the summer) will be implemented in the ECDL. I believe he will give me and Tara the schematic for it at some point and we may build it later. 

I configured the TEC. We have 2 Peltier elements, the Thorlabs TEC3-2.5 and the TEC1.4-6. I used the TEC3-2.5 in the TEC, since its operating current is lower and works better with the controller we bought. I was stupid and burned myself pretty badly on the Peltier element because I had the wrong settings, and will remember that it heats up easily (be CAREFUL before touching). The TEC is working! It can adjust temperature, although the readings it gives is in resistance (which can be converted to temperature using the attached Excel file). I'm going to play with it more tomorrow to see if I can tweak the settings and understand what every button does. 

Today I also figured out how to make the laser diode wire to the 9 pin D-sub connector from the LDX-3412 current source that we found. I built the entire thing in the electrical shop, but haven't put in the reverse bias power supply. I will try to do this tomorrow to see if I can get the laser working. Since the TEC is working, I may try and tune the wavelength of the laser output and see what the noise looks like. 

Attachment 1: thermistor_temperature-resistance.xlsx  246 kB
ELOG V3.1.3-